OpenStudy (anonymous):

power series - radius of convergence I started out with the series: 1+2x+4x^2/(2!)^2+6x^3/(3!)^2+... so I did the ratio test for sum (0,∞) of (2n)!x^n/(n!)^2 and got x lim n->∞ of 2/n+1 The answer is 1/4, but I don't understand how that's the answer. I obviously did something wrong because the radius of convergence I get is 0.. I think

OpenStudy (turingtest):

(2n)!x^n/(n!)^2 ^^^^ this is not the right representation of the sequence of terms

OpenStudy (turingtest):

I still can't get 1/4 though, I must admit...

OpenStudy (anonymous):

It is the right representation--I checked with my professor this afternoon. \[\sum_{0}^{\infty}\frac{ (2n)!x^n }{ (n!)^2 }\] I feel that my problem was I typed the sequence incorrectly (go figure). \[1+2x+\frac{ 4!x^2 }{ (2!)^2 }+\frac{ 6!x^3 }{ (3!)^2 }+\frac{ 8!x^4 }{ (4!)^2 }+\frac{ 10!x^5 }{ (5!)^2 }+...\] I did \[\lim_{n \rightarrow \infty} \frac{ (2(n+1))!x^{n+1} }{ ((n+1)!)^2 }\times \frac{ ((n!)^2 }{ (2n)!x^n }\] and canceled to get \[x \lim_{n \rightarrow \infty} \frac{ 4n+2 }{ n+1 } \approx 4x\] and \[4x < 1\] so \[x < \frac{ 1 }{ 4 }\] and the radius of convergence is \[\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

Thank you for all your help though! I love that someone on here is familiar with calc 2 material.